# S2R-HDR: A Large-Scale Rendered Dataset for HDR Fusion

### v1

@@ 5 Experiments
Table 2: Experimental results on SCT [41] and Challenge123 [21] dataset with ground-truth. We first trained the two baseline networks on the S2R-HDR dataset, followed by simulation-to-real knowledge transfer on the SCT and Challenge123 training sets using the S2R-Adapter. In contrast, the other methods were directly trained on the SCT and Challenge123 training sets. The results marked with * are those recalculated using images provided by  [41]. The best results are in bolded.
Methods Train/Fine-tune/Test on SCT [41] Train/Fine-tune/Test on Challenge123 [21]
PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$ HDR-VDP2 PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$ HDR-VDP2
NHDRRNet [50] 36.68 39.61 0.9590 0.9853 63.72 37.82 26.75 0.9769 0.9632 53.38
DHDRNet [19] 40.05 43.37 0.9794 0.9924 65.50 37.83 29.62 0.9707 0.9705 51.32
AHDRNet [49] 42.08 45.30 0.9837 0.9943 67.30 40.44 28.13 0.9877 0.9703 54.58
DiffHDR [52] 42.77 47.11 0.9854 0.9957 69.43 38.78 26.85 0.9890 0.9745 53.38
HDR-Transformer [27] 42.39 46.35 0.9844 0.9948 67.73 40.70 28.72 0.9881 0.9731 54.63
SCTNet [41] 42.55 47.51 0.9850 0.9952 69.22 40.65 28.73 0.9882 0.9721 54.35
@@
Figure 7: Visual results on the SCT [41] datasets (left) and Challenge123 [21] datasets (right) with ground-truth training data. Our method effectively eliminates artifacts caused by motion occlusions, delivering superior visual quality.
Datasets. In line with latest research [41, 21], we train and evaluate our models on recent HDR datasets, including:
  • SCT Dataset. The SCT dataset [41] includes 108 training samples and 36 test samples, each capturing dynamic scenes with significant foreground or camera motion.
  • Challenge123 Dataset. The Challenge123 dataset [21] is a complex multi-exposure HDR dataset, which was collected using a vivo X90 Pro+, and contains 96 training samples and 27 test samples.
Experiment details. We select the three latest methods [27, 41, 21] as our baselines: HDR-Transformer [27] and SCTNet [41] are transformer-based approaches, while SAFNet [21] is a CNN-based approach. When training these methods on our S2R-HDR dataset, we first generate three different exposure LDR images from the original HDR images. Then, we apply the same data augmentation and training strategy.
Evaluation metrics. We employ commonly used metrics, including PSNR and SSIM, along with HDR-VDP2 [28], a metric specifically designed for HDR evaluation. PSNR and SSIM are computed in both linear and $\mu$-law tone-mapped domains, denoted as $-\ell$ and $-\mu$, respectively.
@@ 5.1 Results
Results on test datasets with ground truth. To validate the effectiveness of our method (S2R-HDR dataset and S2R Adapter), we conducted a comparative study on the latest SCT [41] and Challenge123 [21] datasets against seven widely adopted HDR approaches, including both CNN-based [19, 50, 49, 41], Transformer-based [27, 41] and diffusion-based [52] models. We selected the latest SCTNet and SAFNet as our baseline networks, where SCTNet represents the Transformer-based method and SAFNet represents the CNN-based method. Specifically, we first trained the two baseline networks on the S2R-HDR dataset, followed by synthetic-to-real knowledge transfer on the SCT and Challenge123 training sets using the S2R Adapter. In contrast, the other methods were directly trained on the SCT and Challenge123 training sets.
As shown in Tab. 2, our method achieved the best results on both datasets. In terms of the PSNR-mu metric, our approach demonstrated at least a 0.6dB improvement over both baseline networks on PSNR-$\mu$, and notably achieved a significant 2dB gain on the Challenge123 dataset across both baselines. Additionally, we provide a comparative analysis of visual effects, as illustrated in Fig. 7. Our method effectively reduces artifacts caused by motion occlusions, delivering superior visual quality. We further conduct a visual analysis of the effectiveness of S2R-Adapter on the SCT dataset. For more details, please refer to Sec. A.2.
Table 3: Experimental result on SCT [41] and Challenge123 [21] without ground-truth. We report the testing results of baselines pre-trained on real-world datasets generalizing to the SCT and Challenge123 test datasets, followed by the S2R-Adapter test-time adaptation results of SCTNet and SAFNet pre-trained on S2R-HDR. The best results are in bolded.
Methods Test on SCT [41] Test on Challenge123 [21]
Train PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$ HDR-VDP2 Train PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$ DHR-VDP2
DiffHDR [52] Challenge123 32.33 35.35 0.9497 0.9582 64.16 SCT 34.59 25.33 0.9748 0.9603 52.83
HDR-Transformer [27] 31.94 34.23 0.9518 0.9503 62.70 34.48 24.60 0.9744 0.9573 52.69
SCTNet [41] 32.60 35.93 0.9535 0.9639 63.50 34.57 25.07 0.9753 0.9599 52.09
SAFNet [21] 35.14 38.77 0.9619 0.9868 64.03 34.26 25.50 0.9718 0.9590 52.69
@@
SAFNet w S2R (Ours) 36.28 47.23 0.9586 0.9949 68.40 43.01 32.29 0.9884 0.9831 57.38
@@
Results on test datasets without ground truth. We conduct the following experiment to validate the effectiveness of S2R-Adapter when generalizing to unseen test datasets without ground truth. The pre-trained models are tested on unseen datasets SCT [41] and Challenge123 [21], where no ground-truth label is available for the models, and each test sample is seen only once. As shown in Tab. 3, compared with SCTNet and SAFNet trained on existing real-world datasets, models trained on our S2R-HDR dataset coupled with our S2R-Adapter can more effectively generalize to the unseen target domain. For instance, using SAFNet on the SCT dataset, our approach achieved a 1.1dB improvement in PSNR-$\mu$ and an 8.46dB improvement in $-\ell$ compared to the best baselines. The S2R-Adapter alone provides 1.39dB and 3.38dB improvements in PSNR-$\mu$ and $-\ell$, respectively.
With our test-time adaptation framework, models pre-trained on our dataset can effectively generalize to unseen images. A qualitative comparison using real-captured data without ground truth is illustrated in Fig. 8. Our method effectively alleviates artifacts in highlight areas during nighttime and reduces ghosting caused by large motions. More visual comparisons are available in Sec. A.4.
@@ 5.2 Ablation Study
Effectiveness of S2R-HDR dataset. To evaluate the effectiveness of the S2R-HDR dataset, we select the three latest methods [27, 41, 21] as baselines, which include both transformer-based and CNN-based approaches. Additionally, we chose the two most recent datasets, SCT [41] and Challenge123 [21], as comparative datasets. We train the three baseline methods on the SCT dataset, the Challenge123 dataset, and our S2R-HDR dataset, then evaluate on both the SCT and Challenge123 datasets to assess the generalization capability of each model across datasets. Furthermore, given the domain gap between synthetic datasets (such as S2R-HDR) and real-world datasets (SCT and Challenge123), we also fine-tune the models trained on the synthetic S2R-HDR dataset on the real datasets, following the approach in [29].
As shown in Tab. 4, the model trained on our dataset surprisingly outperforms the one trained directly on Challenge123 when evaluated on the same dataset. Moreover, models trained on either the SCT or Challenge123 datasets suffer significant performance degradation during cross-validation, indicating their limited generalization capability. In contrast, models trained solely on our S2R-HDR dataset—without any exposure to SCT or Challenge123—demonstrate superior cross-dataset generalization, highlighting the high quality and robustness of our dataset. Additionally, models trained on S2R-HDR require only minimal fine-tuning on SCT or Challenge123 to achieve state-of-the-art performance. Across all three tested methods, models trained on S2R-HDR outperformed those trained directly on SCT or Challenge123, achieving at least a 0.4 dB improvement in PSNR-$\mu$. These results confirm the effectiveness of our S2R-HDR dataset in enhancing model robustness and generalization for HDR reconstruction tasks. We further show the visualization results of our S2R-HDR dataset comparison experiments in Sec. A.3.
Table 4: Experimental results on the effectiveness of the S2R-HDR dataset. Models trained on our S2R-HDR dataset demonstrate superior cross-dataset generalization, highlighting the high quality of our dataset. Additionally, only minimal fine-tuning on SCT or Challenge123 is needed to achieve state-of-the-art performance. The results with both the training and test datasets are the same dataset in underline. The best results are in bolded.
Methods Training Testing on SCT [41] Tesing on Challenge123 [21]
PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$ PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$
[…]
@@
Fine-tune 43.25 47.36 0.9877 0.9957 42.40 30.48 0.9912 0.9797
[…]
@@
Fine-tune 43.22 47.28 0.9872 0.9961 42.10 30.18 0.9914 0.9798
[…]
@@
Fine-tune 43.03 48.79 0.9831 0.9958 43.30 31.59 0.9914 0.9819
Table 5: Ablation study of the S2R-Adapter using the SAFNet model [21] pre-trained on the S2R-HDR dataset. The experiments are conducted on the SCT dataset [41]. When both branches work together with learned scale factors $\alpha_{s}$ and $\alpha_{t}$, optimal performance is achieved. In the case of non-learnable $\alpha_{s}$ and $\alpha_{t}$, their values are set to 1.
Baseline Fine-tune Share Transfer Learned PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$
@@
Table 6: Ablation study of the S2R-Adapter Framework under test-time adaptation without GT data. The baseline is the SAFNet [21] pre-trained on our S2R-HDR Dataset. The test data is the SCT Dataset. TS denotes the teach-student framework, Adapter refers to our shared and target branch adapters, and Unc indicates scale factor adjustment with uncertainty.
Baseline TS Adapter Unc PSNR-$\mu$ $-\ell$ SSIM-$\mu$ $-\ell$
@@
Effectiveness of S2R-Adapter’s two branches. To validate the effectiveness of the knowledge-sharing branch and knowledge-transfer branch designed in our Adapter method, we conducted ablation experiments on the SCT dataset using SAFNet as the baseline to evaluate the impact of each branch on the experimental results. As shown in Tab. 5, we tested the effect of using each branch individually. Results indicate that using only the knowledge-sharing branch outperforms simple fine-tuning, suggesting that this branch effectively learns shared knowledge, thereby reducing the forgetting of pre-trained knowledge. Meanwhile, using only the knowledge-transfer branch leads to a more substantial improvement, further confirming the significant differences between synthetic and real data. When both branches work together with learned scale factors $\alpha_{s}$ and $\alpha_{t}$, optimal performance is achieved.
Effectiveness of knowledge control. We conduct experiments to show that our method better facilitates knowledge control than simple fine-tuning, effectively alleviating knowledge forgetting. Specifically, we first train the SAFNet [21] on the S2R-HDR as a pre-trained model. Then, we apply simple fine-tuning and our adapter-based fine-tuning for domain adaptation on the SCT [41] dataset and subsequently test the models on the original S2R-HDR training set to measure knowledge forgetting. As shown in Tab. 7, S2R-Adapter effectively adapts to the SCT dataset while minimizing knowledge forgetting, demonstrating better preservation of pre-trained knowledge.
Table 7: Experiment results on knowledge control using SAFNet [21] on the SCT [41] and S2R-HDR datasets.
SAFNet [21] Test on SCT [41] Test on S2R-HDR
PSNR-$\mu$ $-\ell$ PSNR-$\mu$ $-\ell$
Fine-tune on SCT [41] 43.03 48.79 35.52 29.40
S2R-Adapter on SCT [41] 43.33 48.90 35.95 29.80
Effectiveness of S2R-Adapter framework under test-time adaptation. We validate the S2R-Adapter Framework’s effectiveness during test-time adaptation through ablation studies on the SCT dataset, using SAFNet as the baseline, pre-trained on our S2R-HDR dataset. As shown in Tab. 6, the teacher-student framework enhances results by making the test-time adaptation process more robust. Most improvements are from our shared and transfer branch adapters. Additionally, dynamically adjusting the scale factor between the adapter branches based on uncertainty measurement allows for better control of shared and transferred knowledge across varying domain shifts, further enhancing performance.

### v2

@@ 5 Experiments
Table 2: Experimental results on SCT (Tel et al., 2023) and Challenge123 (Kong et al., 2024) dataset with ground-truth. We first trained the two baseline networks on the S2R-HDR dataset, followed by simulation-to-real domain adaptation on the SCT and Challenge123 training sets using the S2R-Adapter. In contrast, the other methods were directly trained on the SCT and Challenge123 training sets. The results marked with * are those recalculated using images provided by  Tel et al. (2023).
Methods Train/Adaptation/Test on SCT (Tel et al., 2023) Train/Adaptation/Test on Challenge123 (Kong et al., 2024)
$\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$ HDR-VDP2 $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$ HDR-VDP2
NHDRRNet (Yan et al., 2020) 36.68 39.61 0.9590 0.9853 63.72 37.82 26.75 0.9769 0.9632 53.38
DHDRNet (Kalantari and Ramamoorthi, 2017) 40.05 43.37 0.9794 0.9924 65.50 37.83 29.62 0.9707 0.9705 51.32
AHDRNet (Yan et al., 2019) 42.08 45.30 0.9837 0.9943 67.30 40.44 28.13 0.9877 0.9703 54.58
DiffHDR (Yan et al., 2023b) 42.77 47.11 0.9854 0.9957 69.43 38.78 26.85 0.9890 0.9745 53.38
HDR-Transformer (Liu et al., 2022) 42.39 46.35 0.9844 0.9948 67.73 40.70 28.72 0.9881 0.9731 54.63
SCTNet (Tel et al., 2023) 42.55 47.51 0.9850 0.9952 69.22 40.65 28.73 0.9882 0.9721 54.35
@@
Datasets. In line with the latest research (Tel et al., 2023; Kong et al., 2024), we train and evaluate our models on recent HDR datasets: the SCT Dataset(Tel et al., 2023), which contains 108 training samples and 36 test samples featuring dynamic scenes with significant foreground or camera motion; and the Challenge123 Dataset(Kong et al., 2024), a complex multi-exposure HDR dataset collected using a vivo X90 Pro+, comprising 96 training samples and 27 test samples.
Experiment details. We select the three latest methods (Liu et al., 2022; Tel et al., 2023; Kong et al., 2024) as our baselines: HDR-Transformer (Liu et al., 2022) and SCTNet (Tel et al., 2023) are transformer-based approaches, while SAFNet (Kong et al., 2024) is a CNN-based approach. When training these methods on our S2R-HDR dataset, we first generate three different exposure LDR images from the original HDR images. Then, we apply the same data augmentation and training strategy. Regarding adaptation steps, for supervised adaptation with ground-truth labels, we fine-tune the model for about 30 epochs. For the self-supervised mode (test-time adaptation without ground-truth), the model adapts as it processes the data during test time, so each sample is seen only once.
Evaluation metrics. We employ commonly used metrics, including PSNR and SSIM, along with HDR-VDP2 (Mantiuk et al., 2011), a metric designed for HDR evaluation. PSNR and SSIM are computed in both linear and $\mu$-law tone-mapped domains, denoted as -$\ell$ and $\mu$, respectively.
@@ 5.1 Results
Results on test datasets with ground truth. To validate the effectiveness of our method (S2R-HDR dataset and S2R Adapter), we conducted a comparative study on the latest SCT (Tel et al., 2023) and Challenge123 (Kong et al., 2024) datasets against seven widely adopted HDR approaches, including both CNN-based (Kalantari and Ramamoorthi, 2017; Yan et al., 2020; 2019; Tel et al., 2023), Transformer-based (Liu et al., 2022; Tel et al., 2023) and diffusion-based (Yan et al., 2023b) models. We selected the latest SCTNet and SAFNet as our baseline networks, where SCTNet represents the Transformer-based method and SAFNet represents the CNN-based method. Specifically, we first trained the two baseline networks on the S2R-HDR dataset, followed by synthetic-to-real domain adaptation on the SCT and Challenge123 training sets using the S2R Adapter. In contrast, the other methods were directly trained on the SCT and Challenge123 training sets.
As shown in Table 2, our method achieved the best results on both datasets. In terms of the PSNR-$\mu$ metric, our approach demonstrated at least a 0.6dB improvement over both baseline networks on PSNR-$\mu$, and notably achieved a significant 2dB gain on the Challenge123 dataset across both baselines. Additionally, we provide a comparative analysis of visual effects, as illustrated in Figure 6. Our method effectively reduces artifacts caused by motion occlusions, delivering superior visual quality. We further visualize the difference maps of model output on the SCT dataset before and after applying the S2R-Adapter in Appendix A.2. The results show that our S2R-Adapter effectively reduces the domain gap, especially in texture-rich areas.
Figure 6: Visual results on the SCT (Tel et al., 2023) datasets (left) and Challenge123 (Kong et al., 2024) datasets (right) with ground-truth training data. Our method effectively eliminates artifacts caused by motion occlusions, delivering superior visual quality.
Table 3: Experimental result on SCT (Tel et al., 2023) and Challenge123 (Kong et al., 2024) without ground-truth (test-time adaptation). We report the testing results of baselines pre-trained on real-world datasets generalizing to the SCT and Challenge123 test datasets, followed by the S2R-Adapter test-time adaptation results of SCTNet and SAFNet pre-trained on S2R-HDR.
Methods Test on SCT (Tel et al., 2023) Test on Challenge123 (Kong et al., 2024)
Train $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$ HDR-VDP2 Train $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$ DHR-VDP2
DiffHDR (Yan et al., 2023b) Challenge123 32.33 35.35 0.9497 0.9582 64.16 SCT 34.59 25.33 0.9748 0.9603 52.83
HDR-Transformer (Liu et al., 2022) 31.94 34.23 0.9518 0.9503 62.70 34.48 24.60 0.9744 0.9573 52.69
SCTNet (Tel et al., 2023) 32.60 35.93 0.9535 0.9639 63.50 34.57 25.07 0.9753 0.9599 52.09
SAFNet (Kong et al., 2024) 35.14 38.77 0.9619 0.9868 64.03 34.26 25.50 0.9718 0.9590 52.69
@@
SAFNet w S2R-Adapter (Ours) 36.28 47.23 0.9586 0.9949 68.40 43.01 32.29 0.9884 0.9831 57.38
Results on test datasets without ground truth. We conduct the following experiment to validate the effectiveness of S2R-Adapter when generalizing to unseen test datasets where ground truth labels are not available for the adapted models. The pre-trained models are adapted to unseen target datasets SCT (Tel et al., 2023) and Challenge123 (Kong et al., 2024), without accessing the ground truth of the target domain during adaptation. Models will see each test sample only once. As shown in Table 3, compared with SCTNet and SAFNet trained on existing real-world datasets, models trained on our S2R-HDR dataset coupled with our S2R-Adapter can more effectively generalize to the unseen target domain. For instance, using SAFNet on the SCT dataset, our approach achieved a 1.1dB improvement in PSNR-$\mu$ and an 8.46dB improvement in PSNR-$\ell$ compared to the best baselines. The S2R-Adapter alone provides 1.39dB and 3.38dB improvements in PSNR-$\mu$ and PSNR-$\ell$, respectively.
@@
With our test-time adaptation framework, models pre-trained on our dataset can effectively generalize to unseen images. A qualitative comparison using real-captured data without ground truth is illustrated in Figure 7. Our method effectively alleviates artifacts in highlight areas during nighttime and reduces ghosting caused by large motions. More visual comparisons are available in Appendix A.4.
@@ 5.2 Ablation Study
Table 4: Experimental results on the effectiveness of the S2R-HDR dataset. We test the cross-dataset generalization of models trained on different datasets. Models trained on our S2R-HDR dataset achieve superior generalization and require only minimal fine-tuning on SCT or Challenge123 to reach state-of-the-art performance. The best results are in bold.
Methods Training Testing on SCT (Tel et al., 2023) Testing on Challenge123 (Kong et al., 2024)
$\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$ $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
HDR-Transformer (Liu et al., 2022) SCT (Tel et al., 2023) 42.39 46.35 0.9844 0.9948 34.48 24.60 0.9744 0.9573
Challenge123 (Kong et al., 2024) 31.94 34.23 0.9518 0.9503 40.70 28.72 0.9881 0.9731
SCT & Challenge123 42.09 44.64 0.9864 0.9947 39.52 28.23 0.9892 0.9744
@@
S2R-HDR Fine-tune on SCT or Challenge123 43.25 47.36 0.9877 0.9957 42.40 30.48 0.9912 0.9797
SCTNet (Tel et al., 2023) SCT (Tel et al., 2023) 42.55 47.51 0.9850 0.9952 34.57 25.07 0.9753 0.9599
Challenge123 (Kong et al., 2024) 32.60 35.93 0.9535 0.9639 40.65 28.73 0.9882 0.9721
SCT & Challenge123 40.00 42.79 0.9800 0.9935 40.04 28.21 0.9898 0.9750
@@
S2R-HDR Fine-tune on SCT or Challenge123 43.22 47.28 0.9872 0.9961 42.10 30.18 0.9914 0.9798
SAFNet (Kong et al., 2024) SCT (Tel et al., 2023) 42.66 48.38 0.9831 0.9955 34.26 25.50 0.9718 0.9590
Challenge123 (Kong et al., 2024) 35.14 38.77 0.9619 0.9868 41.88 29.73 0.9897 0.9784
SCT & Challenge123 42.12 45.14 0.9853 0.9941 41.61 29.72 0.9901 0.9788
@@
S2R-HDR Fine-tune on SCT or Challenge123 43.03 48.79 0.9831 0.9958 43.30 31.59 0.9914 0.9819
Effectiveness of S2R-HDR dataset. To evaluate the effectiveness of the S2R-HDR dataset, we select the three latest methods (Liu et al., 2022; Tel et al., 2023; Kong et al., 2024) as baselines, which include both transformer-based and CNN-based approaches. Additionally, we chose the two most recent datasets, SCT (Tel et al., 2023) and Challenge123 (Kong et al., 2024), as comparative datasets. We train the three baseline methods on the SCT dataset, the Challenge123 dataset, a merged dataset combining SCT and Challenge123, and our S2R-HDR dataset, then evaluate their generalization by testing on both SCT and Challenge123. Furthermore, given the domain gap between synthetic datasets (such as S2R-HDR) and real-world datasets (SCT and Challenge123), we also fine-tune the models trained on the synthetic S2R-HDR dataset on the real datasets, following the approach in Niklaus et al. (2021).
As shown in Table 4, the model trained on our dataset surprisingly outperforms the one trained directly on Challenge123 when evaluated on the same dataset. Moreover, models trained on either the SCT or Challenge123 datasets suffer significant performance degradation during cross-validation, indicating their limited generalization capability. In contrast, models trained solely on our S2R-HDR dataset—without any exposure to SCT or Challenge123—demonstrate superior cross-dataset generalization, highlighting the high quality and robustness of our dataset. Additionally, models trained on S2R-HDR require only minimal fine-tuning on SCT or Challenge123 to achieve state-of-the-art performance. Across all three tested methods, models trained on S2R-HDR outperformed those trained directly on SCT or Challenge123, achieving at least a 0.4 dB improvement in PSNR-$\mu$. These results confirm the effectiveness of our S2R-HDR dataset in enhancing model robustness and generalization for HDR fusion tasks. We further show the visualization results of our S2R-HDR dataset comparison experiments in Appendix A.5. We also compare our S2R-HDR dataset with the Kalantari (Kalantari and Ramamoorthi, 2017) and Real-HDRV(Deghosting) (Shu et al., 2024) datasets in Appendix A.6.
Table 5: Ablation study of the S2R-Adapter using the SAFNet model (Kong et al., 2024) pre-trained on the S2R-HDR dataset. The experiments are conducted on the SCT dataset (Tel et al., 2023). When both branches work together with learned scale factors $\alpha_{s}$ and $\alpha_{t}$, optimal performance is achieved. In the case of non-learnable $\alpha_{s}$ and $\alpha_{t}$, their values are set to 1.
Baseline Fine-tune Share Transfer Learned $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
@@
Table 6: Ablation study of the S2R-Adapter Framework under test-time adaptation without GT data. The baseline is the SAFNet (Kong et al., 2024) pre-trained on S2R-HDR Dataset. The test data is the SCT Dataset. TS stands for the teacher-student framework, Adapter for shared and target branch adapters, and Unc for scale factor adjustment with uncertainty.
Baseline TS Adapter Unc $\mu PSNR$-$\ell$ $\mu SSIM$-$\ell$
@@
Effectiveness of S2R-Adapter’s two branches. To validate the effectiveness of the knowledge-sharing branch and knowledge-transfer branch designed in our Adapter method, we conducted ablation experiments on the SCT dataset using SAFNet as the baseline to evaluate the impact of each branch on the experimental results. As shown in Table 5.2, we tested the effect of using each branch individually. Results indicate that using only the knowledge-sharing branch outperforms simple fine-tuning, suggesting that this branch effectively learns shared knowledge, thereby reducing the forgetting of pre-trained knowledge. Meanwhile, using only the knowledge-transfer branch leads to a more substantial improvement, further confirming the significant differences between synthetic and real data. When both branches work together with learned scale factors $\alpha_{s}$ and $\alpha_{t}$, optimal performance is achieved.
Effectiveness of knowledge control. We conduct experiments to show that our method better facilitates knowledge control than simple fine-tuning, effectively alleviating knowledge forgetting. Specifically, we first train the SAFNet (Kong et al., 2024) on the S2R-HDR as a pre-trained model. Then, we apply simple fine-tuning and our adapter-based fine-tuning for domain adaptation on the SCT (Tel et al., 2023) dataset and subsequently test the models on the original S2R-HDR training set to measure knowledge forgetting. As shown in Table 7, S2R-Adapter effectively adapts to the SCT dataset while minimizing knowledge forgetting, demonstrating better preservation of pre-trained knowledge.
Table 7: Experiment results on knowledge control using SAFNet (Kong et al., 2024) on the SCT (Tel et al., 2023) and S2R-HDR datasets. The result shows that our S2R-Adapter better alleviates knowledge forgetting.
SAFNet (Kong et al., 2024) Test on SCT (Tel et al., 2023) Test on S2R-HDR
$\mu PSNR$-$\ell$ $\mu PSNR$-$\ell$
Fine-tune on SCT (Tel et al., 2023) 43.03 48.79 35.52 29.40
S2R-Adapter on SCT (Tel et al., 2023) 43.33 48.90 35.95 29.80
Effectiveness of S2R-Adapter framework under test-time adaptation. We validate the S2R-Adapter Framework’s effectiveness during test-time adaptation through ablation studies on the SCT dataset, using SAFNet as the baseline, pre-trained on our S2R-HDR dataset. As shown in Table 5.2, the teacher-student framework enhances results by making the test-time adaptation process more robust. Most improvements are from our shared and transfer branch adapters. Additionally, dynamically adjusting the scale factor between the adapter branches based on uncertainty measurement allows for better control of shared and transferred knowledge across varying domain shifts, further enhancing performance.